# Scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions

### v2

@@ 1 Introduction
$\mathbb{H}\!\!$ $=$ $\!\!-\frac{{\rm i}}{q^{2}-q^{-2}}\bigg{(}\sum_{J=1}^{2L-2}{\cal O}_{J}-(q+q^{-1})\,\big{(}\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{y}\sigma_{2}^{y}+\sigma_{2L-1}^{x}\sigma_{2L}^{x}+\sigma_{2L-1}^{y}\sigma_{2L}^{y}\big{)}-(q^{2}-q^{-2})\,(\sigma^{z}_{2L}-\sigma^{z}_{1})$ (1.3)
-- $\!\!2\,(\sigma^{z}_{1}\sigma^{z}_{2}-\hat{{\bf 1}})+(q^{2}+q^{-2})\ (\sigma^{z}_{2L-1}\sigma^{z}_{2L}-\hat{{\bf 1}})\bigg{)}\ ,$
@@
In this paper we perform a systematic study of the low energy spectrum of the Hamiltonian (1.3) at large system size $L\gg 1$. It is carried out via a mixture of methods, including a numerical analysis of the Bethe Ansatz equations as well as the powerful analytical technique of the ODE/IQFT correspondence. It turns out that the ODEs describing the scaling limit fall within the class of differential equations considered in refs.[15, 26], where the universal behaviour of the vertex model with quasi-periodic boundary conditions imposed was studied. They are described in detail in sec. 3.2 below. As for the Bethe Ansatz equations, they read as
@@
The paper is organized as follows. In section 2 we present a formula for the matrix elements of the Baxter $Q$ operator that was used in our numerical work. It is valid in the sector $S^{z}=0$ and for a one parameter family of open boundary conditions. The outcomes of our study of the low energy spectrum of the Hamiltonian (1.3) in the regime (1.2) is given in section 3. The first subsection thereof, for the most part, is a review of the results contained in refs.[22, 23]. In the next two, the ODE/IQFT correspondence is described and, on the basis of this, the so-called ‘quantization condition’ is obtained. The latter is what allows us to perform a full characterization of the low energy space of states of the lattice model, which is detailed in section 4. In section 5, the formula for the partition function appearing in the scaling limit of the lattice model is given, which may be of interest to those studying non-rational boundary CFT. The last section is devoted to a discussion and includes a summary of the main results.
@@ 2 The $Q$ operator
A method is required in order to extract the Bethe roots corresponding to the state $|\Psi_{L_{\rm in}}\rangle$ which seeds the RG trajectory. Typically for $|\Psi_{L_{\rm in}}\rangle$ generic, $\{\zeta_{m}\}$ are complex numbers that do not resemble a simple pattern in the complex plane, see Fig. 1 for an example. Finding all of the possible solution sets of the Bethe Ansatz equations (1.8) even for $L_{\rm in}\lesssim 10$ is impossible to carry out on a modern laptop because the system is to complicated. Moreover, searching for the set $\{\zeta_{m}\}$ by applying the Newton method to (1.8), trying out various initial approximations, is time consuming and not guaranteed to work.
@@
Below we present the explicit formula for the $Q$-operator that was used in our analysis. It allowed us to go beyond the results of the previous papers [22, 23]. Our formula was obtained based on the results of ref.[28], which is part of an interesting group of recent papers [28, 29, 30, 31] appearing in the mathematics literature. Among other things, they give $\mathbb{Q}(\zeta)$ as a trace of a monodromy matrix over a $q$-oscillator representation for models associated with the rational [29] and trigonometric [30, 31] $R$-matrix for $\mathfrak{sl}(2)$ with a two parameter family of open Boundary Conditions (BCs). The matrix elements of $\mathbb{Q}(\zeta)$ from [28] take the form of an infinite sum, which converges only in a certain parameteric domain that excludes the model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant BCs we are considering. As such, some analysis was required in order to bring the expression to a form, which is literally applicable to the case at hand and is efficient for numerical purposes.
@@
It can be interpreted as an operator $R(\zeta)=R_{I,J}(\zeta)$ acting on the tensor product $\mathbb{C}_{I}^{2}\otimes\mathbb{C}^{2}_{J}$, where the convention is used such that $\mathbb{C}_{J}^{2}$ with $J=0$ stands for the auxiliary space, while for $J=1,\dots,2L$ it is the $J$-th factor of the quantum space
@@
The transfer-matrix, which is graphically depicted in Fig. 2, is given by [^3] The formula for the transfer matrix (2.6) may be equivalently written in index notation: $\big{[}\mathbb{T}(\zeta q^{-1})\big{]}^{b_{1}\dots b_{2L}}_{a_{1}\dots a_{2L}}=q^{-2L}\sum_{c_{1}\dots c_{2L}=\pm}\sum_{\alpha_{1}\dots\alpha_{2L+1}=\pm\atop\beta_{1}\dots\beta_{2L+1}=\pm}$ $\big{[}K^{+}(\zeta)\big{]}^{\beta_{2L+1}}_{\alpha_{2L+1}}\,\big{[}R(\zeta\eta^{-1}_{2L})\big{]}^{\alpha_{2L+1}b_{2L}}_{\alpha_{2L}c_{2L}}\dots\big{[}R(\zeta\eta^{-1}_{1})\big{]}^{\alpha_{2}b_{1}}_{\alpha_{1}c_{1}}\,$ (2.5) $\big{[}K^{-}(\zeta)\big{]}^{\alpha_{1}}_{\beta_{1}}\,\big{[}R(\zeta\eta_{1})\big{]}^{c_{1}\beta_{1}}_{a_{1}\beta_{2}}\dots\big{[}R(\zeta\eta_{2L})\big{]}^{c_{2L}\beta_{2L}}_{a_{2L}\beta_{2L+1}}\,.$ Here $\big{[}K^{\pm}(\zeta)\big{]}^{\beta}_{\alpha}$ with $\alpha,\beta=\pm$ stand for the entries of the diagonal matrices (2.7) and similar for $\big{[}R(\zeta)\big{]}^{\beta b}_{\alpha a}$, e.g., $\big{[}R(\zeta)\big{]}^{+-}_{-+}=q-q^{-1}$ and $\big{[}R(\zeta)\big{]}^{-+}_{+-}=(q-q^{-1})\,\zeta$.
@@
We will use the notation [^4] The transfer-matrix as defined in formula (2.6) for arbitrary inhomogenieties $\eta_{J}$ and $\epsilon=0$ is related to $\tau(u)$ given by (2.9) of ref. [23] via a similarity transformation by a diagonal matrix and an overall multiplicative factor. Namely, $\mathbb{T}^{(0)}(\zeta)=2^{4L}\mbox{e}^{-4Lu-2u}\bigg{(}\prod^{2L}_{J=1}\rho(-u+\tfrac{{\rm i}\gamma}{2}-\delta_{J})\bigg{)}\ \mathcal{U}\,\tau(u)\,\mathcal{U}^{-1}.$ (2.10) Here $\mathcal{U}=G_{1}(\delta_{1})\otimes\dots\otimes G_{2L}(\delta_{2L})$ with $G(u)={\rm{diag}}(1,\mbox{e}^{-u})$, while the parameters $u,\gamma,\delta_{J}$ need to be identified with $\zeta,q,\eta_{J}$ as $\zeta=\mbox{e}^{-2u},\qquad q=\mbox{e}^{{\rm i}\gamma},\qquad\eta_{J}=\mbox{e}^{-2\delta_{J}}\,.$ (2.11) Also, the function $\rho(u)$ is given by $\rho(u)=\frac{1}{2}\big{(}\cos(2\gamma)-\cosh(2u)\big{)}$.
@@
For our purposes it is sufficient to focus on the sector where the eigenvalue of the $z$projection of the total spin operator is zero, i.e., $S^{z}=0$. This is because for the model possessing $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariance, the states come in multiplets $\mathcal{M}_{\mathcal{S}}$ each of which has a representative in that sector. Let the tuples $(a_{1}a_{2}\dots a_{2L})$ and $(b_{1}b_{2}\dots b_{2L})$ with $a_{J},b_{J}=\pm$ be the input/output indices for the space $\mathscr{V}_{2L}$ (2.4). To present the formula for the $Q$-operator, introduce
@@
The matrix elements of $\mathbb{Q}(\zeta)$, valid in the sector $S^{z}=0$, are given by
@@
$\mathbb{Q}(\zeta^{-1})=\zeta^{-2L}\ \mathbb{Q}(\zeta)\,,\qquad\qquad\mathbb{Q}(0)=\mathds{1}\,.$ (2.26)
For arbitrary values of the inhomogeneities $\eta_{J}$ and parameter $\epsilon$, the $Q$-operator obeys the commutativity conditions (2.1) as well as the $TQ$ relation with the transfer-matrix
@@
$+$ $\big{(}\epsilon+q^{-1}\,\zeta\big{)}\,\big{(}1+\zeta\,q^{-1}\,\epsilon\big{)}\,f(q^{+1}\zeta)\,\mathbb{Q}\big{(}\zeta\,q^{-2}\big{)}\qquad\qquad(S^{z}=0)\,,$
where
$f(\zeta)=(1-\zeta^{2})\prod_{J=1}^{2L}\big{(}\zeta-\eta_{J}^{-1}\big{)}\big{(}\zeta-\eta_{J}\big{)}\ .$ (2.28)
The above implies the Bethe Ansatz equations for the integrable model in the usual way: One considers both sides of eq. (2) evaluated on a common eigenvector. In view of (2.26), the eigenvalues of $\mathbb{Q}(\zeta)$ take the form
@@
These are equivalent to the Bethe Ansatz equations obtained in the original work [32], specialized to the sector $S^{z}=0$ and with $\xi^{+}=-\xi^{-}$. The parameters should be identified as $(q,\eta_{J},\epsilon)\mapsto(\mbox{e}^{\eta},\mbox{e}^{-2u_{J}},\mbox{e}^{2\xi^{\pm}})$, while $\zeta_{m}\mapsto\mbox{e}^{-2v_{m}}$.
@@ 3.1 Preliminaries
$E\asymp L\,e_{\infty}+f_{\infty}+\frac{\pi v_{\rm F}}{L}\bigg{(}\frac{p^{2}}{n+2}+\frac{b^{2}}{n}-\frac{1}{12}+{\tt d}\bigg{)}+o(L^{-1-\varepsilon})\ .$ (3.3)
Here the Fermi velocity $v_{\rm F}$ is non-universal and depends on the overall multiplicative normalization of the Hamiltonian. In our conventions for $\mathbb{H}$, one has
@@
$p=\frac{1}{2}\,\big{(}2{\cal S}+1+{\tt w}\,(n+2)\big{)}\,\qquad{\rm with}\qquad{\tt w}=-1\ ,$ (3.5)
where ${\cal S}$ is the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin of $|\Psi_{L}\rangle$, while the notation ${\tt d}$ stands for ‘descendent’ and is a non-negative integer ${\tt d}=0,1,2,\ldots$. Also, the correction term $o(L^{-1-\varepsilon})$ contains an infinitesimally small $\varepsilon>0$.
@@
$\{\zeta_{m}\}_{m=1}^{M}=\{\zeta_{m}^{(+)}\}_{m=1}^{M+\frac{\mathfrak{m}}{2}}\,\cup\,\{\zeta_{m}^{(-)}\}_{m=1}^{M-\frac{\mathfrak{m}}{2}}\qquad\qquad{\rm with}\qquad\qquad\zeta_{m}^{(\pm)}\gtrless 0\qquad\qquad(M=L-{\cal S})\,.$ (3.9)
Comparison of the energy computed from the solution set $\{\zeta_{m}\}$ via eq. (1.9) with that coming from the direct diagonalization of the Hamiltonian, for small lattice sizes, shows that the corresponding Bethe states are low energy states of the spin chain provided $\mathfrak{m}\ll L$. The large $L$ asymptotic of the energy turns out to obey (3.3) with ${\tt d}=0$ so long as there are no significant gaps in the distribution of Bethe roots along the positive and negative rays $\zeta\lessgtr 0$. Computing the value of $b(L)$ from the definition (3.7), (3.6) one finds it to be a real number, which turns out to satisfy the large $L$ asymptotic behaviour [23]
@@
which can be arranged to be an arbitrary real number. Let’s compare the asymptotic formula for the energy (3.3) for the RG trajectory with the general CFT prediction for a lattice system with open boundary conditions imposed [36, 35]:
$E\asymp L\,e_{\infty}+f_{\infty}+\frac{\pi v_{\rm F}}{L}\bigg{(}\Delta-\frac{c}{24}\bigg{)}+o(L^{-1-\varepsilon})\ .$ (3.12)
Here $\Delta$ is the conformal dimension, while $c$ stands for the central charge. One finds that
$\Delta-\frac{c}{24}=\frac{p^{2}}{n+2}+\frac{s^{2}}{n}-\frac{1}{12}+{\tt d}$ (3.13)
with ${\tt d}=0$. Thus we conclude that the spectrum of scaling dimensions develops a continuous component labeled by the parameter $s\in\mathbb{R}$.
@@
$s=\pm{\rm i}\,\Big{(}-p-\frac{1}{2}-a\,\Big{)}\qquad\qquad{\rm with}\qquad\qquad a=0,1,2,\ldots<-p-\tfrac{1}{2}\ \qquad\qquad({\tt d}=0)\,.$ (3.14)

### v3

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
$\mathbb{H}\!\!$ $=$ $\!\!-\frac{{\rm i}}{q^{2}-q^{-2}}\bigg{(}\sum_{J=1}^{2L-2}{\cal O}_{J}-(q+q^{-1})\,\big{(}\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{y}\sigma_{2}^{y}+\sigma_{2L-1}^{x}\sigma_{2L}^{x}+\sigma_{2L-1}^{y}\sigma_{2L}^{y}\big{)}$
$\!\!-(q^{2}-q^{-2})\,(\sigma^{z}_{2L}-\sigma^{z}_{1})-2\,(\sigma^{z}_{1}\sigma^{z}_{2}-\hat{{\bf 1}})+(q^{2}+q^{-2})\ (\sigma^{z}_{2L-1}\sigma^{z}_{2L}-\hat{{\bf 1}})\bigg{)}\ ,$ (1.3)
@@
In this paper we perform a systematic study of the low energy spectrum of the Hamiltonian (1) at large system size $L\gg 1$. It is carried out via a mixture of methods, including a numerical analysis of the Bethe Ansatz equations as well as the powerful analytical technique of the ODE/IQFT correspondence. It turns out that the ODEs describing the scaling limit fall within the class of differential equations considered in refs.[15, 26], where the universal behaviour of the vertex model with quasi-periodic boundary conditions imposed was studied. They are described in detail in sec. 3.2 below. As for the Bethe Ansatz equations, they read as
@@
The paper is organized as follows. In section 2 we present a formula for the matrix elements of the Baxter $Q$ operator that was used in our numerical work. It is valid in the sector $S^{z}=0$ and for a one parameter family of open boundary conditions. The outcomes of our study of the low energy spectrum of the Hamiltonian (1) in the regime (1.2) is given in section 3. The first subsection thereof, for the most part, is a review of the results contained in refs.[22, 23]. In the next two, the ODE/IQFT correspondence is described and, on the basis of this, the so-called ‘quantization condition’ is obtained. The latter is what allows us to perform a full characterization of the low energy space of states of the lattice model, which is detailed in section 4. In section 5, the formula for the partition function appearing in the scaling limit of the lattice model is given, which may be of interest to those studying non-rational boundary CFT. The last section is devoted to a discussion and includes a summary of the main results.
@@ 2 The $Q$ operator
[…]
@@
Below we present the explicit formula for the $Q$ operator that was used in our analysis. It allowed us to go beyond the results of the previous papers [22, 23]. Our formula was obtained based on the results of ref.[28], which is part of an interesting group of recent papers [28, 29, 30, 31] appearing in the mathematics literature. Among other things, they give $\mathbb{Q}(\zeta)$ as a trace of a monodromy matrix over a $q$-oscillator representation for models associated with the rational [29] and trigonometric [30, 31] $R$-matrix for $\mathfrak{sl}(2)$ with a two parameter family of open Boundary Conditions (BCs). The matrix elements of $\mathbb{Q}(\zeta)$ from [28] take the form of an infinite sum, which converges only in a certain parameteric domain that excludes the model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant BCs we are considering. As such, some analysis was required in order to bring the expression to a form, which is literally applicable to the case at hand and is efficient for numerical purposes.
@@
It can be interpreted as an operator $R(\zeta)=R_{I,J}(\zeta)$ acting on the tensor product $\mathbb{C}_{I}^{2}\otimes\mathbb{C}^{2}_{J}$, where $\mathbb{C}_{J}^{2}$ with $J=0$ stands for the auxiliary space, while for $J=1,\dots,2L$ it is the $J$-th factor of the quantum space
@@
The transfer-matrix, which is graphically depicted in Fig. 2, is given by
@@
We will use the notation [^3] The transfer-matrix as defined in formula (2.5) for arbitrary inhomogenieties $\eta_{J}$ and $\epsilon=0$ is related to $\tau(u)$ given by (2.9) of ref. [23] via a similarity transformation by a diagonal matrix and an overall multiplicative factor. Namely, $\mathbb{T}^{(0)}(\zeta)=2^{4L}\mbox{e}^{-4Lu-2u}\bigg{(}\prod^{2L}_{J=1}\rho(-u+\tfrac{{\rm i}\gamma}{2}-\delta_{J})\bigg{)}\ \mathcal{U}\,\tau(u)\,\mathcal{U}^{-1}.$ Here $\mathcal{U}=G_{1}(\delta_{1})\otimes\dots\otimes G_{2L}(\delta_{2L})$ with $G(u)={\rm{diag}}(1,\mbox{e}^{-u})$, while the parameters $u,\gamma,\delta_{J}$ need to be identified with $\zeta,q,\eta_{J}$ as $\zeta=\mbox{e}^{-2u},\qquad q=\mbox{e}^{{\rm i}\gamma},\qquad\eta_{J}=\mbox{e}^{-2\delta_{J}}\,.$ Also, the function $\rho(u)$ is given by $\rho(u)=\frac{1}{2}\big{(}\cos(2\gamma)-\cosh(2u)\big{)}$.
@@
For our purposes it is sufficient to focus on the sector where the eigenvalue of the $z$projection of the total spin operator is zero, i.e., $S^{z}=0$. This is because for the model possessing $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariance, the states come in multiplets $\mathcal{M}_{\mathcal{S}}$ each of which has a representative in that sector. Let the tuples $(a_{1}a_{2}\dots a_{2L})$ and $(b_{1}b_{2}\dots b_{2L})$ with $a_{J},b_{J}=\pm$ be the input/output indices for the space $\mathscr{V}_{2L}$ (2.4). To present the formula for the $Q$ operator, introduce
@@
The matrix elements of $\mathbb{Q}(\zeta)$, valid in the sector $S^{z}=0$, are given by [^4] Written in index notation, the transfer matrix (2.5) takes the form: $\big{[}\mathbb{T}(\zeta q^{-1})\big{]}^{b_{1}\dots b_{2L}}_{a_{1}\dots a_{2L}}=q^{-2L}\sum_{c_{1}\dots c_{2L}=\pm}\sum_{\alpha_{1}\dots\alpha_{2L+1}=\pm\atop\beta_{1}\dots\beta_{2L+1}=\pm}$ $\big{[}K^{+}(\zeta)\big{]}^{\beta_{2L+1}}_{\alpha_{2L+1}}\,\big{[}R(\zeta\eta^{-1}_{2L})\big{]}^{\alpha_{2L+1}b_{2L}}_{\alpha_{2L}c_{2L}}\dots\big{[}R(\zeta\eta^{-1}_{1})\big{]}^{\alpha_{2}b_{1}}_{\alpha_{1}c_{1}}\,$ $\big{[}K^{-}(\zeta)\big{]}^{\alpha_{1}}_{\beta_{1}}\,\big{[}R(\zeta\eta_{1})\big{]}^{c_{1}\beta_{1}}_{a_{1}\beta_{2}}\dots\big{[}R(\zeta\eta_{2L})\big{]}^{c_{2L}\beta_{2L}}_{a_{2L}\beta_{2L+1}}\,.$ Here $\big{[}K^{\pm}(\zeta)\big{]}^{\beta}_{\alpha}$ with $\alpha,\beta=\pm$ stand for the entries of the diagonal matrices (2.6) and similar for $\big{[}R(\zeta)\big{]}^{\beta b}_{\alpha a}$, e.g., $\big{[}R(\zeta)\big{]}^{+-}_{-+}=q-q^{-1}$ and $\big{[}R(\zeta)\big{]}^{-+}_{+-}=(q-q^{-1})\,\zeta$.
@@
$\mathbb{Q}(\zeta^{-1})=\zeta^{-2L}\ \mathbb{Q}(\zeta)\,,\qquad\qquad\mathbb{Q}(0)={\bf 1}\,.$ (2.23)
The following comment is in order regarding the relation of the $Q$ operator whose matrix elements are given by (2.20) with that studied in ref. [28]. In fact, in the latter work two $Q$ operators $\mathbb{Q}^{(a)}$ with $a=1,2$ are introduced. In order to specialize to the one parameter family of open boundary conditions being considered here one should restrict the parameters ϵ¯±subscript¯italic-$\bar{\epsilon}_{\pm}$ and ϵ±subscriptitalic-$\epsilon_{\pm}$ in that paper as $\epsilon_{+}/\epsilon_{-}=\bar{\epsilon}_{+}/\bar{\epsilon}_{-}=\epsilon$. Then, choosing a representation for the $q$ oscillator algebra, formula (5.13) from [28] provides an expression for the matrix elements of the $Q$ operators in terms of an infinite sum $\sum_{m\geq 0}g(m)$ For the sector $S^{z}=0$ and generic values of the parameters this sum diverges for both $\mathbb{Q}^{(1)}$ and $\mathbb{Q}^{(2)}$ so that eq. (5.13) becomes inapplicable. This is because the summand $g(m)$ tends to a finite, nonvanishing limit as $m\to\infty$. Our formula (2.20) essentially coincides with $\lim_{m\to\infty}g(m)$. The limiting value is the same whether or not we started from $\mathbb{Q}^{(1)}$ or $\mathbb{Q}^{(2)}$ so one may only obtain a single $Q$ operator in this way. Remarkably, we have checked for small lattice sizes $L=2,3,4,\ldots$ that $\mathbb{Q}(\zeta)$ from (2.20) obeys the commutativity conditions (2.1) as well as the $TQ$ relation with the transfer-matrix
@@
$+$ $\big{(}\epsilon+q^{-1}\,\zeta\big{)}\,\big{(}1+\zeta\,q^{-1}\,\epsilon\big{)}\,f(q^{+1}\zeta)\ \mathbb{Q}\big{(}\zeta\,q^{-2}\big{)}\qquad\qquad(S^{z}=0)\,.$
Here
$f(\zeta)=(1-\zeta^{2})\prod_{J=1}^{2L}\big{(}\zeta-\eta_{J}^{-1}\big{)}\big{(}\zeta-\eta_{J}\big{)}\ ,$ (2.25)
while the values of the inhomogeneities $\eta_{J}$ and parameter $\epsilon$ are assumed to be generic. The Bethe Ansatz equations for the integrable model follow as usual: One considers both sides of eq. (2) evaluated on a common eigenvector. In view of (2.23), the eigenvalues of $\mathbb{Q}(\zeta)$ take the form
@@
These are equivalent to the Bethe Ansatz equations obtained in the original work [32], specialized to the sector $S^{z}=0$ and with $\xi^{+}=-\xi^{-}$. Also, the parameters should be identified as $(q,\eta_{J},\epsilon)\mapsto(\mbox{e}^{\eta},\mbox{e}^{-2u_{J}},\mbox{e}^{2\xi^{\pm}})$, while $\zeta_{m}\mapsto\mbox{e}^{-2v_{m}}$. We find it surprising that a formula like (2.20) exists and its further exploration may be worthwhile. At the same time, since our work is focused on the study of the scaling limit of an integrable lattice system, we believe that it is not the place to do this here.
@@ 3.1 Preliminaries
$E\asymp L\,e_{\infty}+f_{\infty}+\frac{\pi v_{\rm F}}{L}\bigg{(}\frac{p^{2}}{n+2}+\frac{b^{2}}{n}-\frac{1}{12}+\texttt{d}\bigg{)}+o(L^{-1-\varepsilon})\ .$ (3.3)
The Fermi velocity $v_{\rm F}$ is non-universal and depends on the overall multiplicative normalization of the Hamiltonian. In our conventions for $\mathbb{H}$, one has
@@
$p=\frac{1}{2}\,\big{(}2{\cal S}+1+\texttt{w}\,(n+2)\big{)}\,\qquad{\rm with}\qquad\texttt{w}=-1\ ,$ (3.5)
where ${\cal S}$ is the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin of $|\Psi_{L}\rangle$, while the notation d stands for ‘descendent’ and is a non-negative integer $\texttt{d}=0,1,2,\ldots$. Also, the correction term $o(L^{-1-\varepsilon})$ contains an infinitesimally small $\varepsilon>0$.
@@
[…]
@@
which can be arranged to be an arbitrary real number. Here and below we use the symbol $\operatorname*{slim}$ for ‘scaling limit’ to emphasize that there is additional non-trivial input involved in taking the number of sites to infinity. Let’s compare the asymptotic formula for the energy (3.3) for the RG trajectory with the general CFT prediction for a lattice system with open boundary conditions imposed [36, 35]:
$E\asymp L\,e_{\infty}+f_{\infty}+\frac{\pi v_{\rm F}}{L}\bigg{(}\Delta-\frac{c}{24}\bigg{)}+o(L^{-1-\varepsilon})\ ,$ (3.12)
where $\Delta$ is the conformal dimension, while $c$ stands for the central charge. One finds that
$\Delta-\frac{c}{24}=\frac{p^{2}}{n+2}+\frac{s^{2}}{n}-\frac{1}{12}+\texttt{d}$ (3.13)
with $\texttt{d}=0$. Thus we conclude that the spectrum of scaling dimensions develops a continuous component labeled by the parameter $s\in\mathbb{R}$.
@@
$s=\pm{\rm i}\,\Big{(}-p-\frac{1}{2}-a\,\Big{)}\qquad\qquad{\rm with}\qquad\qquad a=0,1,2,\ldots<-p-\tfrac{1}{2}\ \qquad\qquad(\texttt{d}=0)\,.$ (3.14)